How many grams of sodium azide are required to produce 28.0 g of nitrogen?
1
Expert's answer
2020-03-26T03:02:35-0400

Solution:

The decomposition of sodium azide (NaN3) is:

2NaN=> 2Na(s) + 3N2(g)

According to the chemical equation: n(NaN3)/2 = n(Na)/2 = n(N2)/3.

n(NaN3)/2 = n(N2)/3;

3*n(NaN3) = 2*n(N2).


a)

m(NaN3) = ??? g;

M(NaN3) = Ar(Na) + 3*Ar(N) = 23 + 3*14 = 65 (g/mol);

n(NaN3) = m(NaN3) / M(NaN3);

n(NaN3) = m(NaN3) / 65

b)

m(N2) = 28.0 g;

M(N2) = 2*Ar(N) = 2*14 = 28 (g/mol);

n(N2) = m(N2) / M(N2);

n(N2) = 28.0 / 28 = 1.0 mol


Consequently,

3*n(NaN3) = 2*n(N2);

n(NaN3) = 2*n(N2) / 3

m(NaN3) / 65 = (2 * 1.0) / 3

m(NaN3) / 65 = 0.667;

m(NaN3) = 65*0.667 = 43.35

m(NaN3) = 43.35 g


Answer: 43.35 g of sodium azide are required to produce 28.0 g of nitrogen.